# Hybrid-Task Meta-Learning: A Graph Neural Network Approach for Scalable and Transferable Bandwidth Allocation

### v2

@@ IV-C2 Unsupervised Learning
IV-C3 Computational Complexity
We compare the computational complexity of our GNN with the iterative algorithm introduced in Section IV-A. In cellular systems, both algorithms will be implemented in each transmission time interval with a duration of less than 1 ms. Thus, we are interested in the inference complexity of our GNN, i.e., the number of operations to be executed to obtain the bandwidth allocation in each transmission time interval.
Inference complexity of our GNN
To compute the embedding of each vertex, we need to compute the output of the FNN in Fig. 1. We denote the number of layers of the FNN by $L_{\mathrm{FNN}}$ and the number of neurons in the $\ell$-th layer by $m_{\mathrm{FNN}}^{\ell}$. Then, the number of multiplications required to compute the output of the $\ell$-th layer is $m_{\mathrm{FNN}}^{\ell}\cdot m_{\mathrm{FNN}}^{\ell+1}$ and the total number of multiplications for computing the embedding is $M_{\mathrm{FNN}}=\sum\nolimits_{\ell=1}^{L_{\mathrm{FNN}}}m_{\mathrm{FNN}}^{\ell}\cdot m_{\mathrm{FNN}}^{\ell+1}$ [10]. After obtaining the embeddings of $K$ users, the number of multiplications required by $f_{\mathtt{Softmax}}(\bm{x})$ and $f_{\mathtt{Readout}}(\bm{y})$ is $2K$. Therefore, the inference complexity of the GNN-based bandwidth allocation policy is
$O_{\mathrm{GNN}}=O(K\cdot(M_{\mathrm{FNN}}+2)).$ (16)
Complexity of the iterative algorithm
In each iteration of the optimization algorithm, we assign a small portion of the normalized surplus bandwidth, denoted by $\Delta\tilde{w}$, to a user that can maximize the objective function. The algorithm needs to compute the objective function $K$ times and find the best user. We denote the complexity for computing the objective function by $\Omega$, then the complexity of the iterative algorithm is given by
$O_{\mathrm{iter}}=O\left(K\cdot\frac{w_{\mathrm{S}}}{\Delta w}\cdot\Omega\right),$ (17)
where $w_{\mathrm{S}}/{\Delta{w}}$ represents the number of iterations used in the iterative algorithm.
Complexity comparison
To obtain bandwidth allocation in each transmission time interval, the transmitter either uses the forward propagation algorithm to compute the outcome of the GNN or executes the iterative algorithm. From eqs. (16) and (17), we can see that the computational complexity of our GNN and the iterative algorithm increase linearly with the number of users. Recall that $M_{\mathrm{FNN}}$ in eq. (16) is quite limited. In contrast, the complexity of the iterative algorithm also increases with the amount of surplus bandwidth and the resource block and thus depends on the channels of the users. In addition, the computing complexity for evaluating the objective function, denoted by $\Omega$ in eq. (17), in each iteration of the optimization algorithm could also be extremely high. Thus, the inference complexity of the GNN is much lower than the complexity of the iterative optimization algorithm.
@@ IV-D3 Fine-Tuning Set and Evaluation Set in Meta-Testing
1 Initialize the number of training epochs, $N$, and the learning rate of the target testing task, $\beta_{\theta}$.
2 Select the $i$-th task from the fine-tuning set and the evaluation set: $\mathcal{T}_{i}^{\mathrm{F}}\in\mathcal{T}^{\mathrm{F}}$ and $\mathcal{T}_{i}^{\mathrm{E}}\in\mathcal{T}^{\mathrm{E}}$.
@@ 5       Randomly select $J$ samples from task $\mathcal{T}_{i}^{\mathrm{F}}$
6       Calculate loss of in the fine-tuning set according to (15), denoted by $f^{\mathrm{L,F}}(\theta^{n})$.
@@
10Randomly select $J^{\prime}$ samples from task $\mathcal{T}_{i}^{\mathrm{E}}$.
@@
1 Randomly initialize the training parameters for all the tasks, $\phi$, the number of meta-training epochs, $M$, the learning rate of meta-training, $\beta_{\phi}$, and the learning rate of each task, $\beta_{\theta}$.
@@
8                   Calculate the loss function in the support set according to (15), denoted by ${f}^{\mathrm{L,S}}(\theta_{i}^{m,n})$.  Update the parameters by: $\theta_{i}^{m,n+1}=\theta_{i}^{m,n}+{\beta_{\theta}}\nabla_{\theta}{f}^{\mathrm{L,S}}(\theta_{i}^{m,n}).$
@@
10            Select a batch of $I^{\prime}$ tasks from the query set: $\mathcal{T}_{i^{\prime}}^{\mathrm{Q}}\in\mathcal{T}^{\mathrm{Q}},\quad i^{\prime}\in\{1,2,\cdots,I^{\prime}\}$. for $i^{\prime}=1,\cdots,I^{\prime}$ do
@@
16      Calculate the loss function in meta-training: ${f}^{\mathrm{L,Meta},m}\left(\phi^{m}\right)=\frac{1}{I}\sum\limits_{i=1}^{I}{f}_{i}^{\mathrm{L,Q}}\left(\theta_{i}^{m,N}\right)$.  Update the initial parameters: $\phi^{m+1}=\phi^{m}+\beta_{\phi}\nabla_{\phi}{f}^{\mathrm{L,Meta},m}\left(\phi^{m}\right)$.
17
@@
Algorithm 4 Meta-Training of Hybrid-Task Meta-Learning
@@ IV-D4 Meta-Training of Proposed HML Algorithm
Fig. 2(b) illustrates the tasks and tasksets used in the meta-training and meta-testing of the proposed HML algorithm. The difference between MAML and HML lies in the selection of tasks from the query set. In MAML, the tasks selected from the query set are identical to those selected from the support set in each meta-training epoch. To improve the generalization ability, in HML, we select different tasks from the query set to train the initial parameters of the GNN. Specifically, $I^{\prime}$ tasks are randomly selected from the query set to estimate the average loss of the GNN parameterized by $\phi^{m}$ in the $m$-th epoch of meta-training. The step-by-step algorithm for meta-training of the proposed HML algorithm is described in Algorithm 4, and the meta-testing algorithm of HML is the same as that of MAML in Algorithm 3.
@@
Single-sided noise spectral density, $N_{0}$174 dBm/Hz
Channel coherence time, $T_{\mathrm{c}}$ $1$ms [28]
Duration of one time slot, $T_{\mathrm{s}}$ $0.125$ms
@@
Parameters $\mathcal{T}^{\mathrm{S}}$ $\mathcal{T}^{\mathrm{Q}}$ $\mathcal{T}^{\mathrm{F}}$ $\mathcal{T}^{\mathrm{E}}$
@@
Shadowing: $p_{u}^{\mathrm{S}}(\psi)=\frac{10/\ln 10}{\sqrt{2\pi}\sigma_{\psi_{\mathrm{dB}}}\psi}\exp\left(-\frac{(10\log_{10}\psi-\mu_{\psi_{\mathrm{dB}}})^{2}}{2\sigma_{\psi_{\mathrm{dB}}}^{2}}\right)$ $\psi_{\mathrm{dB}}\in\{3,4,5\}$ $\psi_{\mathrm{dB}}=8$
Small-scale channels: $p_{u}^{\mathrm{I}}(z|s,\sigma)=\frac{z}{\sigma^{2}}\exp\left(-\frac{z^{2}+s^{2}}{2\sigma^{2}}\right)\cdot I_{0}\left(\frac{zs}{\sigma^{2}}\right)$, $p_{u}^{\mathrm{N}}(z|m,\sigma)=\frac{2m^{m}z^{2m-1}}{\Gamma(m){(2\sigma^{2})}^{m}}\exp\left(-\frac{mz^{2}}{2\sigma^{2}}\right)$, $p_{u}^{\mathrm{R}}(z|\sigma)=\frac{z}{\sigma^{2}}\exp\left(-\frac{z^{2}}{2\sigma^{2}}\right)$ $p_{u}^{\mathrm{I}}(z|s,\sigma),s\in\{1\cdots 5\}$, $p_{u}^{\mathrm{N}}(z|m,\sigma),m\in\{2,\cdots,6\}$ $p_{u}^{\mathrm{R}}(z|\sigma)$
QoS Rewards with different QoS requirements $\max\limits_{\bm{w}}\sum\limits_{u\in\mathcal{U}_{\tau}^{S,\mathcal{I}}}r_{u}^{S,\mathcal{I}}$ $\max\limits_{\bm{w}}\sum\limits_{u\in\mathcal{U}_{\tau}^{\Phi,\xi}}r_{u}^{\Phi,\xi}$, $\Phi\in\{D,E,S\}$, $\xi\in\{\mathcal{I,F}\}$
Values of QoS constraints (Mbps) $r_{\tau}^{S,\mathcal{I}}\in\{1,\cdots,10\}$ $r_{\tau}^{\Phi,\xi}=10$
@@ V-A System Setup
We consider a BS, located at $(0,0)$ m, serving multiple users randomly distributed in a rectangular area, where the coordinates of the users are denoted by $(c_{x,u},c_{y,u})$, where $c_{x,u}$ and $c_{y,u}$ $\in[-100,100]$. When the QoS requirement is secrecy rate, an eavesdropper is randomly located in the above rectangular area. The transmitted signal of each user is a complex Gaussian process with zero-mean and equal variance, $\sigma^{2}=1$. Channel models include large-scale channels and small-scale channels. Specifically, the large-scale channels depend on path loss and shadowing fading, whilst small-scale channels follow Rice, Nakagami, and Rayleigh distributions with various parameters in Table III. The number of neurons in each layer of the GNN is $2/32/64/64/32/1$. Unless otherwise mentioned, the simulation parameters are summarized in Table II, and the parameters of tasksets are defined in Table III.
@@ V-B Performance of GNN
Figure 3: Training losses with different numbers of users, where the secrecy rate in the long blocklength regime is considered, $r_{\tau}^{S,\mathcal{I}}=10$ Mbps, and $W_{\tau}^{S,\mathcal{I}}=100$ MHz.
@@
After the training stage of the unsupervised learning algorithm, we select $1000$ samples from the evaluation set of the same task to evaluate the constraint and reward achieved by the GNN in Fig. 4. The results in Fig. 4(a) show that the secrecy rates of all the scheduled users are equal to or higher than the requirement, $r_{\tau}^{S,\mathcal{I}}=10$ Mbps. The results in Fig. 4(b) show that the sum secrecy rate achieved by the GNN is close to that achieved by the iterative optimization algorithm in Section IV-A (with legend “Optimal”). In other words, the unsupervised learning algorithm can obtain a near-optimal solution.
@@ V-C Meta-Testing Performance of HML
In this subsection, we evaluate the generalization ability of the proposed HML algorithm. The differences between tasks in meta-training and meta-testing are shown in Table. III. In meta-testing, we first select an unseen task that is not included in meta-training. In each training epoch of the meta-testing, $32$ samples are randomly selected from $\mathcal{T}^{\mathrm{F}}$ to fine-tune the GNN, whilst all the $1000$ testing samples from the same task in $\mathcal{T}^{\mathrm{E}}$ are used to evaluate the performance.
@@ V-C1 Different Wireless Channels and QoS Requirements
In this part, we set $W_{\tau}^{S,\mathcal{I}}=100$ MHz and $r_{\tau}^{S,\mathcal{I}}=10$ Mbps for all types of services. The other parameters follow the rules in $\mathcal{T}^{\mathrm{S}}$ and $\mathcal{T}^{\mathrm{Q}}$ as shown in Table. III. We compare the initial performance and sample efficiency of HML with four benchmarks: 1) Optimal, 2) Model-agnostic meta-learning (MAML), 3) Multi-task learning-based transfer learning (MTL Transfer), and 4) Random initialization.
  • Optimal: The optimal solution is obtained by the iterative algorithm detailed in Section IV-A, and its optimality has been proved in [10].
@@
  • MTL Transfer: Transfer learning improves the sample efficiency by fine-tuning the parameters of the pre-trained GNN in a task with fewer training samples. With multi-task learning (MTL), the initial performance is much better than random initialization as the GNN is pre-trained in multiple tasks [37, 43]. To execute MTL transfer learning, we only need to replace the initialization in line  2 of Algorithm 2 by the pre-trained parameters.
@@
In Fig. 5, we consider the average sum of secrecy rates and illustrate the impacts of the number of users, channel models, and coding blocklength on the initial performance and sample efficiency of different methods. The results in Fig. 5 show that HML achieves the best initial average sum secrecy rate and the highest sample efficiency compared with all the benchmarks. In Fig. 5(a), HML can converge in $8$ training epochs. Both MAML and MTL transfer learning takes more than $30$ epochs to converge. Thus, HML can reduce the convergence time by up to $73\%$. After the fine-tuning, the gap between learning methods and the optimal solution is around $1.45$%. In Fig. 5(b), the coding blocklength in meta-testing is also different from that in meta-training. As a result, the gap between the initial performance of HML and the optimal solution is $7.93\%$. After fine-tuning, the gap reduced to $3.74\%$, which is larger than the gap in Fig. 5(a), where the blocklength is the same in meta-training and meta-testing.
@@
Fig. 6 shows the average sum of data rates achieved by different methods. The results indicate that when the reward function and the QoS constraint in meta-testing are different from that in meta-training, the gaps between the initial performance of HML and the optimal solution increase to $13.77\%$ and $14.93\%$ in long and short blocklength regimes, respectively. After fine-tuning, the gaps between the learning methods and the optimal solution are smaller than that in Fig. 5. This is because Shannon’s capacity/achievable rate are two special cases of the secrecy rate in the long/short blocklength regimes when the wiretapped channels are in deep fading. It is easier to learn a good policy when the problem becomes less complicated.

### v3

@@ IV-D3 Fine-Tuning Set and Evaluation Set in Meta-Testing
1Initialize the number of training epochs, NN, and the learning rate of the target testing task, $\beta_{\theta}$.
2
3Select the $i$-th task from the fine-tuning set and the evaluation set: $\mathcal{T}_{i}^{\mathrm{F}}\in\mathcal{T}^{\mathrm{F}}$ and $\mathcal{T}_{i}^{\mathrm{E}}\in\mathcal{T}^{\mathrm{E}}$.
@@ 6  Randomly select $J$ samples from task $\mathcal{T}_{i}^{\mathrm{F}}$
7  Calculate loss in the fine-tuning set according to (16), denoted by $f^{\mathrm{L,F}}(\theta^{n})$.
@@
11Randomly select $J^{\prime}$ samples from task $\mathcal{T}_{i}^{\mathrm{E}}$.
@@
1Randomly initialize the training parameters for all the tasks, $\phi$, the number of meta-training epochs, $M$, the learning rate of meta-training, $\beta_{\phi}$, and the learning rate of each task, $\beta_{\theta}$.
@@
8
9       Calculate the loss function in the support set according to (16), denoted by ${f}^{\mathrm{L,S}}(\theta_{i}^{m,n})$
10       Update the parameters by: $\theta_{i}^{m,n+1}=\theta_{i}^{m,n}+{\beta_{\theta}}\nabla_{\theta}{f}^{\mathrm{L,S}}(\theta_{i}^{m,n}).$
@@
12
13    Select a batch of $I^{\prime}$ tasks from the query set: $\mathcal{T}_{i^{\prime}}^{\mathrm{Q}}\in\mathcal{T}^{\mathrm{Q}},\quad i^{\prime}\in\{1,2,\cdots,I^{\prime}\}$.
14    for $i^{\prime}=1,\cdots,I^{\prime}$ do
@@
20 Calculate the loss function in meta-training: ${f}^{\mathrm{L,Meta},m}\left(\phi^{m}\right)=\frac{1}{I}\sum\limits_{i=1}^{I}{f}_{i}^{\mathrm{L,Q}}\left(\theta_{i}^{m,N}\right)$
21 Update the initial parameters: $\phi^{m+1}=\phi^{m}+\beta_{\phi}\nabla_{\phi}{f}^{\mathrm{L,Meta},m}\left(\phi^{m}\right)$.
22
@@
Algorithm 5 Meta-Training of Proposed HML
@@ IV-D4 Meta-Training of Proposed HML Algorithm
Fig. 2(b) illustrates the tasks and tasksets used in the meta-training and meta-testing of the proposed HML algorithm. The difference between MAML and HML lies in the selection of tasks from the query set. In MAML, the tasks selected from the query set are identical to those selected from the support set in each meta-training epoch. To improve the generalization ability, in HML, we select different tasks from the query set to train the initial parameters of the GNN. Specifically, $I^{\prime}$ tasks are randomly selected from the query set to estimate the average loss of the GNN parameterized by $\phi^{m}$ in the $m$-th epoch of meta-training. In the meta-training stage, our GNN learns the hidden features across the diverse tasks, thus facilitating the adaptation ability in novel tasks in the meta-testing stage. The step-by-step algorithm for meta-training of the proposed HML algorithm is described in Algorithm 5, and the meta-testing algorithm of HML is the same as that of MAML in Algorithm 4.
IV-E Complexity Analysis
IV-E1 Computational Complexity Derivations
We compare the computational complexity of our GNN and HML algorithms with the iterative algorithm given in Algorithm 2. In our considered cellular systems, all the algorithms will be implemented in each transmission time interval with a duration of less than $1$ ms. Thus, we are interested in the inference complexity of the proposed GNN and HML algorithms, i.e., the number of operations to be executed to obtain the bandwidth allocation in each transmission time interval.
To compute the embedding of each vertex in the GNN, we first calculate the output of the FNN depicted in Fig. 1(b). We denote the number of layers of the FNN by $L_{\mathrm{FNN}}$ and the number of neurons in the $\ell$-th layer by $m_{\mathrm{FNN}}^{\ell}$. Then, the number of multiplications required to compute the output of the $\ell$-th layer is $m_{\mathrm{FNN}}^{\ell}\cdot m_{\mathrm{FNN}}^{\ell+1}$, and the total number of multiplications for computing the embedding is $M_{\mathrm{FNN}}=\sum\nolimits_{\ell=1}^{L_{\mathrm{FNN}}}m_{\mathrm{FNN}}^{\ell}\cdot m_{\mathrm{FNN}}^{\ell+1}$ [10]. After obtaining the embeddings of the $K$ scheduled users with sufficiently strong channels, the number of multiplications required by $f_{\texttt{Softmax}}(\bm{x})$ and $f_{\texttt{Readout}}(\bm{y})$ is $2K$. Therefore, the inference complexity of the GNN-based bandwidth allocation policy is
$O_{\mathrm{GNN}}=O(\Omega_{\mathrm{US}}+K\cdot(M_{\mathrm{FNN}}+2)),$ (17)
where $\Omega_{\mathrm{US}}$ represents the computational complexity for the user scheduling process. Since HML employs the same neural network architecture as our GNN, the inference latency of HML is equivalent to that of the GNN and is given by
$O_{\mathrm{HML}}=O_{\mathrm{GNN}}.$ (18)
For the iterative algorithm, as shown in line 2 of Algorithm 2, in each iteration, we assign a small portion of the normalized surplus bandwidth, denoted by $\Delta\tilde{w}$, to a user that can maximize the objective function. The algorithm needs to compute the objective function $K$ times to find the best user in each iteration. We denote the complexity for computing the objective function by $\Omega^{\Phi,\xi}$ , then the complexity of the iterative algorithm is given by
$O_{\mathrm{iter}}=O\left(\Omega_{\mathrm{US}}+K\cdot\frac{w_{\tau,\mathrm{S}}^{\Phi,\xi}}{\Delta w}\cdot\Omega^{\Phi,\xi}\right),$} (19)
where $w_{\tau,\mathrm{S}}^{\Phi,\xi}$ is the surplus bandwidth defined in eq. (12), and $w_{\tau,\mathrm{S}}^{\Phi,\xi}/{\Delta{w}}$ represents the number of iterations used in the iterative algorithm.
IV-E2 Complexity Comparative Analysis
To obtain bandwidth allocation in each transmission time interval, the transmitter either uses the forward propagation algorithm to compute the outcome of the GNN/HML or executes the iterative algorithm. From eqs. (17) and (19), we can see that the computational complexity of our GNN/HML and the iterative algorithm increase linearly with the number of scheduled users, $K$. Recall that $M_{\mathrm{FNN}}$ in eq. (17) is a small number [10], in contrast, although the complexity of the iterative algorithm also depends on the channels of the scheduled users, it increases with the amount of surplus bandwidth, $w_{\mathrm{S}}$, and decreases with the resource block, $\Delta w$. More importantly, the computing complexity for evaluating the objective function, $\Omega^{\Phi,\xi}$, in each iteration of the optimization algorithm could also be extremely high. Therefore, the inference complexity of the GNN is much lower than the complexity of the iterative optimization algorithm.
@@ IV-E2 Complexity Comparative Analysis
Single-sided noise spectral density, $N_{0}$ -174 dBm/Hz
Channel coherence time, $T_{\mathrm{c}}$ 11ms [29]
Duration of one time slot, $T_{\mathrm{s}}$ 0.1250.125ms
@@
Parameters $\mathcal{T}^{\mathrm{S}}$ & $\mathcal{T}^{\mathrm{Q}}$ $\mathcal{T}^{\mathrm{F}}$ & $\mathcal{T}^{\mathrm{E}}$
@@
Shadowing:
$p_{u}^{\mathrm{S}}(\psi)=\frac{10/\ln 10}{\sqrt{2\pi}\sigma_{\psi_{\mathrm{dB}}}\psi}\exp\left(-\frac{(10\log_{10}\psi-\mu_{\psi_{\mathrm{dB}}})^{2}}{2\sigma_{\psi_{\mathrm{dB}}}^{2}}\right)$
$\psi_{\mathrm{dB}}\in\{3,4,5\}$
$\psi_{\mathrm{dB}}=8$
Small-scale channels:
$p_{u}^{\mathrm{I}}(z|s,\sigma)=\frac{z}{\sigma^{2}}\exp\left(-\frac{z^{2}+s^{2}}{2\sigma^{2}}\right)\cdot I_{0}\left(\frac{zs}{\sigma^{2}}\right)$,
$p_{u}^{\mathrm{N}}(z|m,\sigma)=\frac{2m^{m}z^{2m-1}}{\Gamma(m){(2\sigma^{2})}^{m}}\exp\left(-\frac{mz^{2}}{2\sigma^{2}}\right)$,
$p_{u}^{\mathrm{R}}(z|\sigma)=\frac{z}{\sigma^{2}}\exp\left(-\frac{z^{2}}{2\sigma^{2}}\right)$
$p_{u}^{\mathrm{I}}(z|s,\sigma),s\in\{1\cdots 5\}$,
$p_{u}^{\mathrm{N}}(z|m,\sigma),m\in\{2,\cdots,6\}$
$p_{u}^{\mathrm{R}}(z|\sigma)$
QoS Rewards with different QoS requirements $\max\limits_{\bm{w}}\sum\limits_{u\in\mathcal{U}_{\tau}^{S,\mathcal{I}}}r_{u}^{S,\mathcal{I}}$
$\max\limits_{\bm{w}}\sum\limits_{u\in\mathcal{U}_{\tau}^{\Phi,\xi}}r_{u}^{\Phi,\xi}$,
$\Phi\in\{D,E,S\}$, $\xi\in\{\mathcal{I,F}\}$
Values of QoS constraints (Mbps) $r_{\tau}^{S,\mathcal{I}}\in\{1,\cdots,10\}$
$r_{\tau}^{\Phi,\xi}=10$
@@ V-A System Setup
We consider a BS, located at $(0,0)$ m, serving multiple users randomly distributed in a rectangular area, where the coordinates of the users are denoted by $(c_{x,u},c_{y,u})$, where $c_{x,u}$ and $c_{y,u}$ $\in[-100,100]$. When the QoS requirement is secrecy rate, an eavesdropper is randomly located in the above rectangular area. The transmitted signal of each user is a complex Gaussian process with zero-mean and equal variance, $\sigma^{2}=1$. Channel models include large-scale channels and small-scale channels. Specifically, the large-scale channels depend on path loss and shadowing fading, whilst small-scale channels follow Rice, Nakagami, and Rayleigh distributions with various parameters in Table III. The number of neurons in each layer of the GNN is $2/32/64/32/1$. Unless otherwise mentioned, the simulation parameters are summarized in Table II, and the parameters of tasksets are defined in Table III. In real-world communication systems, the BSs capture the information changes in the environment according to the 5G New Radio standard [44].
@@ V-B Performance of GNN
Figure 3: Average training losses with different numbers of users, where the secrecy rate in the long blocklength regime is considered, $r_{\tau}^{S,\mathcal{I}}=10$ Mbps, and $W_{\tau}^{S,\mathcal{I}}=100$ MHz.
@@
After the training stage of the unsupervised learning algorithm, we select $1000$ samples from the evaluation set of the same task to evaluate the constraint and reward achieved by the GNN in Fig. 4. The results in Fig. 4(a) show that the secrecy rates of all the scheduled users are equal to or higher than the requirement, $r_{\tau}^{S,\mathcal{I}}=10$ Mbps. The results in Fig. 4(b) show that the sum secrecy rate achieved by the GNN is close to that achieved by the iterative optimization algorithm in Section IV-A (with legend “Iterative”). In other words, the unsupervised learning algorithm can obtain a near-optimal solution.
@@ V-C Meta-Testing Performance of HML
In this subsection, we evaluate the generalization ability of the proposed HML algorithm. The differences between tasks in meta-training and meta-testing are shown in Table. III. In meta-training, we randomly select the parameters included in $\mathcal{T}^{\mathrm{S}}$ and $\mathcal{T}^{\mathrm{S}}$ in each training epoch. In meta-testing, we first select an unseen task that is not included in meta-training. In each training epoch of the meta-testing, $32$ samples are randomly selected from $\mathcal{T}^{\mathrm{F}}$ to fine-tune the GNN, whilst all the $1000$ testing samples from the same task in $\mathcal{T}^{\mathrm{E}}$ are used to evaluate the performance.
@@ V-C1 Different Wireless Channels and QoS Requirements
In this part, we set $W_{\tau}^{S,\mathcal{I}}=100$ MHz and $r_{\tau}^{S,\mathcal{I}}=10$ Mbps for all types of services. The other parameters follow the rules in $\mathcal{T}^{\mathrm{S}}$ and $\mathcal{T}^{\mathrm{Q}}$ as shown in Table. III. We compare the initial performance and sample efficiency of HML with four benchmarks: 1) Iterative, 2) Model-agnostic meta-learning (MAML), 3) Multi-task learning-based transfer learning (MTL Transfer), and 4) Random initialization.
  • Iterative: The optimal solution is obtained by the iterative algorithm detailed in Section IV-A.
@@
  • MTL Transfer: Transfer learning improves the sample efficiency by fine-tuning the parameters of the pre-trained GNN in a task with fewer training samples. With multi-task learning (MTL), the initial performance is much better than random initialization as the GNN is pre-trained in multiple tasks [38, 45]. To execute MTL transfer learning, we only need to replace the initialization in line  3 of Algorithm 3 by the pre-trained parameters.
@@
In Fig. 5, we consider the average sum of secrecy rates and illustrate the impacts of the number of users, channel models, and coding blocklength on the initial performance and sample efficiency of different methods. The results in Fig. 5 show that HML achieves the best initial average sum secrecy rate and the highest sample efficiency compared with all the benchmarks. In Fig. 5(a), HML can converge in $8$ training epochs. Both MAML and MTL transfer learning take more than $30$ epochs to converge. Thus, HML can reduce the convergence time by up to $73\%$. After the fine-tuning, the gap between learning methods and the optimal solution is around 1.451.45%. In Fig. 5(b), the coding blocklength in meta-testing is also different from that in meta-training. As a result, the gap between the initial performance of HML and the optimal solution is $7.93\%$. After fine-tuning, the gap reduced to $3.74\%$, which is larger than the gap in Fig. 5(a), where the blocklength is the same in meta-training and meta-testing.
@@
Fig. 6 shows the average sum of data rates achieved by different methods. The results indicate that when the reward function and the QoS constraint in meta-testing are different from that in meta-training, the gaps between the initial performance of HML and the optimal solution increase to $13.77\%$ and $14.93\%$ in long and short blocklength regimes, respectively. After fine-tuning, the gaps between the learning methods and the optimal solution are smaller than that in Fig. 5. This is because Shannon’s capacity/achievable rate are two special cases of the secrecy rate in the long/short blocklength regimes when the eavesdropped channels are in deep fading. It is easier to learn a good policy when the problem becomes less complicated.